Consider the following program:

int main()             int abc(int m)           float xyz(float v, int k)
  {                      {                        {
  int a = 3, b;          int i, j = 1;            int i;
  float x = 2.0, y;                               float w = 1.0;

  b = abc(a);            for(i=1; i<=m; i++)      for(i=1; i<=k; i++, w *= v);
  y = xyz(x, a);            j = j * i;
  return 0;
  }                      return j;                return (w);
                         }                        }

When the program is completed, the value of b is __________ and the value of y is ________.

(Go through step by step)
Answer #1

b = 1 * 2 * 3 = 6

y = 2 * 2 * 2 = 8

When the program is completed, the value of b is 6 and the value of y is 8.